# Agile Temporal Discretization for Symbolic Optimal Control

### v1

@@ IV NUMERICAL EXPERIMENTS
In this section we illustrate the benefits of our approach on two different optimal control problems under reach-avoid specifications $\Sigma_{1}^{\text{Reach}}$.
Before stating the numerical results, we provide an explicit way to construct the abstractions (see Remark 1). In this work we construct our abstractions $\mathcal{S}_{2}$ in the same fashion as in [18], except that we allow for mutliple timesteps. To construct the abstract system $\mathcal{S}_{2}=(\mathcal{X}_{2},\mathcal{U}_{2},F_{2})$, we consider a uniform partition of the concrete state space $\mathcal{X}_{1}$. Formally, the abstract state space is defined as
$\mathcal{X}_{2}:=\{q_{1},\dots,q_{N_{q}}\},$ (16)
where $K$ is the number of partition cells. The relation $R\subseteq\mathcal{X}_{1}\times\mathcal{X}_{2}$ is defined such that $(x_{1},q)\in R$ if and only if $x_{1}\in\mathcal{X}_{1}$ lies in the cell represented by $q\in\mathcal{X}_{2}$. We also consider a finite number of inputs $\mathcal{U}_{2}^{\prime}:=\{u_{1},\dots,u_{N_{u}}\}$ and times $\mathcal{T}:=\{\tau_{1},\dots,\tau_{N_{\tau}}\}$ (where we used the same notation as in (7)). To compute a transition function $F_{2}$ that satisfies (8), we use a growth bound [18, Definition VIII.2] to over-approximate the concrete reachable sets $F_{1}(R^{-1}(q),(u,\tau))$ for all $q\in\mathcal{X}_{2}$ and $(u,\tau)\in\mathcal{U}_{2}$. By Theorem 1, this construction ensures that $\mathcal{S}_{1}\preceq_{R}^{\operatorname{FRR}}\mathcal{S}_{2}$.
The abstract reach-avoid specification $\Sigma_{2}^{\text{Reach}}$ is constructed analogously to [18], ensuring that $\Sigma_{1}^{\text{Reach}}\subseteq R^{-1}(\Sigma_{2}^{\text{Reach}})$. Then, by Proposition 1, any controller $\mathcal{C}_{2}$ solving $(\mathcal{S}_{2},\Sigma_{2}^{\text{Reach}})$ can be concretized into a controller $\mathcal{C}_{1}=\mathcal{C}_{2}\circ R$ that solves $(\mathcal{S}_{1},\Sigma_{1}^{\text{Reach}})$.
@@ IV-A Effort-Optimal Control of A Simple Pendulum
To illustrate the theoretical framework of the previous section, we consider a frictionless simple pendulum. This continuous time system is given by $\mathcal{D}=(\mathcal{X},\mathcal{U},f)$ , with $\mathcal{X}=[-\pi,2\pi]$, $\mathcal{U}=[-6,6]$ and dynamics defined by
$f((\theta,\omega),u)=\begin{pmatrix}\omega\\ -\frac{g}{l}\sin{(\theta)+u}\end{pmatrix},$
where $g=9.81[m/s^{2}]$ is gravity, $l$ is the length of the rod and $u$ is the input torque.
Given the transition system $\mathcal{S}_{1}=(\mathcal{X}_{1},\mathcal{U}_{1},F_{1})$ associated with $\mathcal{D}$ as defined in 6, we define the concrete specification $\Sigma_{1}^{\text{Reach}}=[\mathcal{I}_{1},\mathcal{F}_{1},\mathcal{O}_{1}]$ with $\mathcal{I}_{1}=[-\frac{5\pi}{180},\frac{5\pi}{180}]\times[-0.5,0.5]$, $\mathcal{F}_{1}=[\pi-\frac{15\pi}{180},\pi+\frac{15\pi}{180}]\times[-0.5,0.5]$ and $\mathcal{O}_{1}=\varnothing$.
We consider a cost function $c_{1}(x_{1},(u,\tau)):=u^{2}\tau+\epsilon$ [^1] The term $\epsilon$ is added to meet the assumptions of Proposition 2, as without this, $u=0$ leads to a zero cost. In practice, this $\epsilon$ can be chosen arbitrarily close to zero.. One can interpret this cost as a measure of the effort that is needed to reach the specification. The abstract transition cost is then given by $c_{2}(q,(u,\tau)):=u^{2}\tau+\epsilon$, which satisfies condition 14, since $c_{1}$ is independent of the concrete state $x_{1}$.
As a result, by Theorem 2, given a Bellman value function $v_{2}$ and its associated controller $\mathcal{C}_{2}$ for $\mathcal{S}_{2}$, one can construct a suboptimal value function $v_{1}=v_{2}\circ R$ for $\mathcal{S}_{1}$, which provides an upper bound on the worst-case cost under the concretized controller $\mathcal{C}_{1}=\mathcal{C}_{2}\circ R$.
Results are summarized in Table I. We can observe that with our method, we get a feasible abstract problem, which none of the considered timesteps could provide on their own. An example trajectory as well as a representation of the suboptimal value function is given in Figure 2.
Figure 2: Controlled trajectory of the simple pendulum. The red region denotes the target set $\mathcal{T}_{1}$. The heatmap shows the computed suboptimal value function $v_{1}$ for $\mathcal{S}_{1}$. Arrow colors indicate the timestep length used along the closed-loop trajectory—the lighter, the shorter.
Time-step(s) [s] Worst-case effort guarantee ($\mathcal{L}(\mathcal{C}_{1})$)
@@ 0.5 Infeasible
0.1, 0.2, 0.3, 0.4, 0.5 40.959
TABLE I: Worst-case time guarantees for different individual sampling steps. The last row shows the result of our method combining multiple time steps.
IV-B Time-Optimal Path-Planning
We consider a path-planning problem in which a vehicle must navigate through a two-dimensional maze. The system has three state variables: the position $(x,y)$ in the plane and the orientation $\theta$. It is controlled via two inputs: the rear-wheel speed $u_{1}$ and the steering angle $u_{2}$. This model has been previously studied in the context of abstraction-based control [18], and is adapted from Chapter 2.4 of [21].
The continuous-time system is given by $\mathcal{D}=(\mathcal{X},\mathcal{U},f)$, with state space $\mathcal{X}=[-0.1,10.1]\times[-0.1,10.1]\times[-\pi-0.4,\pi+0.4]$, input space $\mathcal{U}\subseteq[-1,1]\times[-1,1]$, and dynamics defined by
$f((x,y,\theta),(u_{1},u_{2}))=\begin{pmatrix}u_{1}\cos(\alpha+\theta)\cos(\alpha)^{-1}\\ u_{1}\sin(\alpha+\theta)\cos(\alpha)^{-1}\\ u_{1}\tan(u_{2})\end{pmatrix},$
where $\alpha=\arctan\left(\frac{\tan(u_{2})}{2}\right)$.
Given the transition system $\mathcal{S}_{1}=(\mathcal{X}_{1},\mathcal{U}_{1},F_{1})$ associated with $\mathcal{D}$ as defined in (6), we define the concrete specification $\Sigma_{1}^{\text{Reach}}=[\mathcal{I}_{1},\mathcal{F}_{1},\mathcal{O}_{1}]$, which constrains only the first two state variables, as illustrated in Figure 3 and its caption.
In the following, we consider the concrete transition cost $c_{1}(x_{1},(u,\tau)):=\tau$, corresponding to the objective of minimizing the total time required to satisfy the reach-avoid specification. The abstract transition cost is defined as $c_{2}(q,(u,\tau)):=\tau$, which satisfies condition (14), since $c_{1}$ is independent of the concrete state $x_{1}$.
We solve the optimal control problem for several time steps and compare the results with our method in Table II.
Time-step(s) [s] Worst-case time guarantee ($\mathcal{L}(\mathcal{C}_{1})$) [s]
0.2 Infeasible
0.3 138.6
0.4 Infeasible
0.5 Infeasible
0.2, 0.3, 0.4, 0.5 97.2
TABLE II: Worst-case time guarantees for different individual sampling steps. The last row shows the result of our method combining multiple time steps.
The benefit of the method is twofold:
  1. 1.
    No need to fine-tune the timestep to determine the precise value that makes the abstract problem feasible.
  2. 2.
    We obtain much better guarantees on the worst-case trajectory.
Figure 3: An example trajectory, projected onto the first two components of the system, for the path-planning problem. The initial set $\mathcal{I}_{1}=\{x_{0}\}$, target set $\mathcal{F}_{1}$, and obstacle set $\mathcal{O}_{1}$ are shown in green, red, and black, respectively. The color of the arrow indicates the used timestep. We see that long time-steps are preferred when available and short steps are used for tight manoeuvers or for readjustment.
See on Figure 3 an example trajectory obtained with our approach.
@@ IV-C Data-driven approach
Growth-bound functions provide over-approximations of the reachable sets $F_{1}(R^{-1}(q),(u,\tau))$, a key requirement for ensuring that Theorem 1 holds. However, growth bound functions become more and more conservative as the timestep increases [18]. This can lead to infeasibility of the abstract problem due to non-determinism, and prevents us from experimenting with longer time steps, as the over-approximation may intersect obstacles and be discarded.
@@
This has the effect of under-approximating the reachable set $F_{1}(R^{-1}(q),(u,\tau))$, as we may overlook a trajectory. As a result, Theorem 1 no longer applies. Nonetheless, it avoids the inclusion of 1-step spurious transitions, as all transitions added in the abstraction are a result of an observed one in the concrete system.
Using this data-driven approach, we are able to consider larger time steps, as shown in Table III, which leads to improved expected worst-case cost. In Figure 4, we show a sample trajectory obtained with this method, starting from the same initial state as in Figure 3. As in the growth-bound-based approach, the controller uses a combination of several different time steps to satisfy the specification.
Time-steps [s] Worst-case time ($\mathcal{L}(\mathcal{C}_{1})$) [s]
0.3, 0.7, 1.1, 1.5, 1.9 94.1
TABLE III: Worst-case time guarantee under different sampling steps for the data-driven approach. We obtain a better guarantee compared to the approach with growth-bound functions, see Table II.
Figure 4: An example trajectory for the path-planning problem for the data driven approach. We can observe the same phenomena than on Figure 3.
@@ V CONCLUSIONS AND FUTURE WORKS
In this work, we introduced a temporal abstraction framework for abstraction-based control of continuous-time systems. By allowing multiple time steps within a unified abstraction, we addressed the trade-off between conservatism and feasibility that arises in fixed-step discretizations. We proved that these abstractions preserve the feedback refinement relation, enabling formal guarantees to be carried over from the abstract to the concrete system. We further demonstrated the effectiveness of our approach through numerical experiments, showing improved worst-case performance and increased flexibility.

### v2

@@ IV NUMERICAL EXPERIMENTS
In this section, we illustrate the benefits of our approach on two different optimal control problems.
Before stating the numerical results, we outline the abstraction construction. We construct our abstractions $\mathcal{S}_{2}$ in the same fashion as in [17], but allow multiple timesteps. To construct the abstract system $\mathcal{S}_{2}=(\mathcal{X}_{2},\mathcal{U}_{2},F_{2})$, we consider a uniform partition of the concrete state space $\mathcal{X}_{1}$. Formally, the abstract state space is defined as $\mathcal{X}_{2}:=\{q_{1},\dots,q_{N_{q}}\}$, where $N_{q}$ is the number of partition cells. The relation $R\subseteq\mathcal{X}_{1}\times\mathcal{X}_{2}$ is defined such that $(x_{1},q)\in R$ if and only if $x_{1}\in\mathcal{X}_{1}$ lies in the cell represented by $q\in\mathcal{X}_{2}$. The transition function $F_{2}$ is computed via a growth bound [17, Definition VIII.2] to over-approximate the concrete reachable sets $F_{1}(R^{-1}(q),(u,\tau))$ for all $q\in\mathcal{X}_{2}$ and $(u,\tau)\in\mathcal{U}_{2}$. By Theorem 1, this ensures that $\mathcal{S}_{1}\preceq_{R}^{\operatorname{FRR}}\mathcal{S}_{2}$. The abstract reach-avoid specification $\Sigma_{2}^{\text{Reach}}$ is constructed analogously to [17], ensuring that $\Sigma_{1}^{\text{Reach}}\subseteq R^{-1}(\Sigma_{2}^{\text{Reach}})$. Then, by Proposition 1, any controller $\mathcal{C}_{2}$ solving $(\mathcal{S}_{2},\Sigma_{2}^{\text{Reach}})$ can be concretized into a controller $\mathcal{C}_{1}=\mathcal{C}_{2}\circ R$ that solves $(\mathcal{S}_{1},\Sigma_{1}^{\text{Reach}})$. Finally, to ensure a fair comparison, we keep abstractions of the same size by fixing the number of input-timesteps pairs.
@@ IV-A Effort-optimal control of a simple pendulum
To illustrate the theoretical framework of the previous section, we consider a frictionless simple pendulum. This continuous time system is given by $\mathcal{D}=(\mathcal{X},\mathcal{U},f)$ , with $\mathcal{X}=[-\pi,2\pi]$, $\mathcal{U}=[-6,6]$ and dynamics defined by $f((\theta,\omega),u)=\begin{pmatrix}\omega\\ -\frac{g}{l}\sin{(\theta)+u}\end{pmatrix}$, where $g=9.81[m/s^{2}]$ is gravity, $l$ is the length of the rod and $u$ is the input torque.
Given the transition system $\mathcal{S}_{1}=(\mathcal{X}_{1},\mathcal{U}_{1},F_{1})$ associated with $\mathcal{D}$ as in Definition 7, we define the concrete specification $\Sigma_{1}^{\text{Reach}}=[\mathcal{I}_{1},\mathcal{F}_{1},\mathcal{O}_{1}]$ with $\mathcal{I}_{1}=[-\frac{5\pi}{180},\frac{5\pi}{180}]\times[-0.5,0.5]$, $\mathcal{F}_{1}=[\pi-\frac{17.5}{180},\pi+\frac{17.5\pi}{180}]\times[-1,1]$ and $\mathcal{O}_{1}=\varnothing$.
We consider a cost function $c_{1}(x_{1},(u,\tau)):=u^{2}\tau+\epsilon$ [^1] The term $\epsilon$ is added to meet the assumptions of Proposition 2, as without this, $u=0$ leads to a zero cost. In practice, this $\epsilon$ can be chosen arbitrarily close to zero.. One can interpret this cost as a measure of the effort that is needed to reach the specification. The abstract transition cost is then given by $c_{2}(q,(u,\tau)):=u^{2}\tau+\epsilon$, which satisfies condition (14), since $c_{1}$ is independent of the concrete state $x_{1}$. As a result, by Theorem 2, given a Bellman value function $v_{2}$ and its associated controller $\mathcal{C}_{2}$ for $\mathcal{S}_{2}$, one can construct a suboptimal value function $v_{1}=v_{2}\circ R$ for $\mathcal{S}_{1}$, which provides an upper bound on the worst-case cost under the concretized controller $\mathcal{C}_{1}=\mathcal{C}_{2}\circ R$.
Results are summarized in Table I. We can observe that with our method we get a feasible abstract problem, which outperforms controllers synthesized with a single timestep. An example trajectory as well as a representation of the suboptimal value function is given in Figure 2.
Figure 2: Controlled trajectory of the simple pendulum. The red region denotes the target set $\mathcal{F}_{1}$. The heatmap shows the computed suboptimal value function $v_{1}$ for $\mathcal{S}_{1}$. Arrow colors indicate the timestep length used along the closed-loop trajectory—the lighter, the shorter.
Timestep(s) [s] Worst-case effort guarantee ($\mathcal{L}(\mathcal{C}_{1})$)
0.1 Infeasible
0.15 81.47
@@ 0.3 Infeasible
0.1, 0.15, 0.2, 25, 0.3 65.70
TABLE I: Worst-case effort guarantees for the simple pendulum problem for different individual sampling steps. The last row shows the results of our method combining multiple timesteps, with a coarser input space discretization.
IV-B Time-optimal path-planning
We consider a path-planning problem in which a vehicle must navigate through a two-dimensional maze. The system has three state variables: the position $(x,y)$ in the plane and the orientation $\theta$. It is controlled via two inputs: wheel speed $u_{1}$ and steering angle $u_{2}$. This model has previously been studied in the context of abstraction-based control [17].
The continuous-time system is given by $\mathcal{D}=(\mathcal{X},\mathcal{U},f)$, with state space $\mathcal{X}=[-0.1,10.1]\times[-0.1,10.1]\times[-\pi-0.4,\pi+0.4]$, input space $\mathcal{U}\subseteq[-1,1]\times[-1,1]$, and dynamics defined by $f((x,y,\theta),(u_{1},u_{2}))=\begin{pmatrix}u_{1}\cos(\alpha+\theta)\cos(\alpha)^{-1}\\ u_{1}\sin(\alpha+\theta)\cos(\alpha)^{-1}\\ u_{1}\tan(u_{2})\end{pmatrix}$, where $\alpha=\arctan\left(\frac{\tan(u_{2})}{2}\right)$.
Given the transition system $\mathcal{S}_{1}=(\mathcal{X}_{1},\mathcal{U}_{1},F_{1})$ associated with $\mathcal{D}$, we define the concrete specification $\Sigma_{1}^{\text{Reach}}$, which constrains only the first two state variables, as illustrated in Figure 3 and its caption.
We consider the concrete transition cost $c_{1}(x_{1},(u,\tau)):=\tau$, which corresponds to minimizing the total time required to satisfy the reach-avoid specification. The abstract transition cost is $c_{2}(q,(u,\tau)):=\tau$, which satisfies condition (14), since $c_{1}$ is independent of the concrete state $x_{1}$.
Timestep(s) [s] Worst-case time guarantee ($\mathcal{L}(\mathcal{C}_{1})$) [s]
@@ 0.5 Infeasible
0.8 144.0
0.2, 0.5, 0.8 114.60
0.3, 0.7, 1.1, 1.5, 1.9 94.1
TABLE II: Worst-case time guarantees for the path planning problem, for different individual sampling steps. The last 2 rows shows the result of our method combining multiple timesteps, with a coarser input space discretization. The last row is obtained with the data-driven approach (Subsection IV-C.
Results are summarized in Table II. Again, a combination of timesteps is used and the worst-case cost is improved. See on Figure 3 an example trajectory obtained with our approach.
Figure 3: An example trajectory, projected onto the first two components of the system, for the path-planning problem. The initial set $\mathcal{I}_{1}=\{x_{0}\}$, target set $\mathcal{F}_{1}$, and obstacle set $\mathcal{O}_{1}$ are shown in green, red, and black, respectively. The color of the arrow indicates the used timestep.
@@ IV-C Data-driven approach
Growth-bound functions provide over-approximations of the reachable sets $F_{1}(R^{-1}(q),(u,\tau))$, necessary so that Theorem 1 holds. However, they become more and more conservative as the timestep increases [17]. This can lead to infeasibility of the abstract problem by introducing non-determinism, and prevents us from experimenting with longer timesteps, as the over-approximation may intersect obstacles and be discarded.
@@
This has the effect of under-approximating the reachable set, as we may overlook a trajectory. As a result, Theorem 1 no longer applies. Using this data-driven approach, we are able to consider larger timesteps, as shown in Table II, which leads to improved expected worst-case cost. In Figure 4, we show a sample trajectory obtained with this method. As with the growth-bound approach, the controller uses a combination of different timesteps.
Figure 4: An example trajectory for the path-planning problem for the data driven approach. The color of the arrows indicate the duration of the timestep.
@@ V CONCLUSIONS AND FUTURE WORKS
In this work, we introduced a temporal abstraction framework for abstraction-based control of continuous-time systems. By allowing multiple timesteps within a unified abstraction, we addressed the trade-off between conservatism and feasibility that arises in fixed-step discretizations. We proved that these abstractions preserve the feedback refinement relation, enabling formal guarantees to be carried over from the abstract to the concrete system. We further demonstrated the effectiveness of our approach through numerical experiments, showing improved worst-case performance without increasing the size of the abstractions.